What is partition function for n dimensional quantum harmonic oscillator.
1
Expert's answer
2020-10-19T13:24:57-0400

1-d quantum oscillator is given by a hamiltonian:

"\widehat H_1 = h\omega_1(\widehat a^+a+\frac1 2)"

n-d quantum oscillator :

"\widehat H_N = \sum_{k=1}^{N}h\omega_1(\widehat a_k^+a_k+\frac1 2)"

The partition function of such model is:

"Z_N(\beta) = Tr_{\mathscr{H}}(e^{-\beta\widehat H_N})"

there is "\mathscr{H} -" hilbert space of n - d quantum oscillator

Rewrite:

"Tr_{\mathscr{H}}(e^{-\beta\widehat H_N}) = \prod^{N}_{k=0}\sum^{\infty}_{n_k=0} e^{-\beta h \omega_k(n_k+\frac 1 2)} ="

"\prod^{N}_{k=0} e^ {\frac {-\beta h \omega_k}{2}} \frac{1}{1-e^ {{-\beta h \omega_k}}} = \frac {2^N}{\prod^{N}_{k=0} sinh(\frac {\beta h \omega_k}{2})}"